Good Evening Tutor,
I am having a hard time with this problem and was hoping to get a bit of explanation to help me out.
Solve the following quadratic equation by factoring or any other technique acceptable.
x^2 - 9x - 4 = 6
Any help on this problem is appreciated. Thank you for your help.

Answer by oberobic(2304)

First, get the equation into standard form: x^2 - 9x - 10 = 0.
Second, think about what factors you could have that would result in -10 and -9.
Well, 10 = 2*5 or 1*10. Subtracting 1 from 10 gives 9, so let's plug that into a couple of factors to see if that works:
(x - 10)(x + 1) = x^2 + x - 10x - 10 = x^2 - 9x - 10.
Answer.